# Model-free source seeking of exponentially convergent unicycle: theoretical and robotic experimental results

### v1

@@ 4.2 Simulation Results
In this subsection, we compare the traditional ESC design commonly found in the literature that is based on first-order Lie bracket design (Dürr et al. (2013)) with the proposed exponentially convergent unicycle ESC design developed for a fourth-order objective function based on third-order Lie bracket. The objective function used in this study is defined as
@@
Figure 3 illustrates the time histories of $x$ and $y$ as well as the planar trajectory of the unicycle. The results clearly demonstrate that the proposed exponentially convergent ESC achieves significantly faster convergence toward the desired equilibrium compared with the traditional ESC method from literature based on first-order Lie bracket design (Dürr et al. (2013)). Specifically, for the fourth-order objective function considered, the traditional ESC fails to converge within the 100 s simulation window, whereas the proposed design successfully converges within approximately 20 s.
@@
Figure 3: Simulation results comparing the proposed exponentially convergent unicycle ESC with the traditional ESC based on first-order Lie bracket design from Dürr et al. (2013) for a fourth-order objective function: (a) $x$ position, (b) $y$ position, and (c) planar trajectory of the proposed unicycle design.
@@ 4.3 Experimental Results
We note here that our design parameters meet the condition of Theorem 1 (see Remark 1). The experimental results are presented in Figures 4 and 5. As shown in Figure 4, the traditional ESC method based on first-order Lie bracket from Dürr et al. (2013) fails to converge to the true minimum of the objective function even after 1200 s, instead settling near (x,y)=(0.98,1.93)(x,y)=(0.98,-1.93).
In contrast, the exponentially convergent unicycle ESC design converges much faster, first reaching the minimum at approximately 350 s and remaining about the true extremum for the duration of the experiment. The new ESC design clearly outperforms the classic literature ESC design and is validated by its ability to reach the minimum in significantly less time than the traditional approach. It is also important to note that a fourth-order objective function exhibits relatively flat regions near the minimum compared to a second-order objective function. The exponentially convergent unicycle ESC design can successfully steer the robot through these flat regions, as evident in Figure 6. Moreover, the variations in the objective function values become very small when the robot is close to the extremum point, confirming that steady-state convergence has been achieved. It is also important to highlight that the proposed design performed well even with expected resolution issues from the motion capturing system feedback due to the very small values of the fourth-order objective function, especially near the extremum. This is a positive indication about the robustness of the proposed design. The reader is directed to watch the experiment in our YouTube channel (Palanikumar et al. (2025a)).
Figure 4: Experimental results for the traditional ESC design based on first-order Lie bracket from Dürr et al. (2013) with a fourth-order objective function. (a) $x$ position (b) $y$ position (c) planar trajectory of the robot. Note that the x-position converges to $0.98$ meters (black-dashed), resulting in a 0.02 meter error, and the y-position converges to 1.93-1.93 (black-dashed), which results in a 0.07 meter error. The true minimum is in red.
Figure 5: Experimental results for the exponentially convergent unicycle ESC design with a fourth-order objective function. (a) $x$ position, (b) $y$ position, (c) planar trajectory of robot, (d) objective function, $J$, (e) linear velocity, $\mathrm{v}$. The reader can watch the experiment at our YouTube channel (Palanikumar et al. (2025a)).
@@
The results of this experiment are presented in Figure 7. As shown, the robot successfully reaches and oscillates around the location of maximum light intensity. The light sensor measurements decrease as the robot approaches the source, and the mean value of the signal is minimized toward the end of the experiment, despite small oscillations caused by the robot’s continuous motion. These results further confirm that the proposed unicycle ESC design can autonomously steer the robot through an unknown and spatially varying light field, thereby validating its real-time, model-free source-seeking capability. The reader is directed to watch this experiment video in YouTube (Palanikumar et al. (2025b)).
@@
Figure 7: Experimental results for light source-seeking with the exponentially convergent ESC model. (a) $x$ position, (b) $y$ position, (c) planar trajectory of robot, (d) objective function, $J$, (e) linear velocity, $\mathrm{v}$. The reader can watch the experiment at our YouTube channel (Palanikumar et al. (2025b)).
@@ 5 Conclusion and future work
This paper provided a novel, first-of-its-kind exponentially convergent unicycle design inspired by significant recent results in higher-order Lie bracket approximations (Grushkovskaya and Eisa (2025); Pokhrel and Eisa (2023)). We proved exponential stability of the proposed design for objective functions that behave locally as fourth-degree polynomial. Additionally, we validated the results by simulations, and more importantly, via experiments. In our experiments, we did not only validate the ability of the proposed design to operate with known objective functions, but we also validated its ability to operate in a complete model-free condition in a light source-seeking experiment.

### v2

@@ 4.2 Simulation Results
In this subsection, we compare the traditional ESC design commonly found in the literature that is based on first-order Lie bracket design (durr2013lie) with the proposed exponentially convergent unicycle ESC design developed for a fourth-order objective function based on third-order Lie bracket. The objective function used in this study is defined as
@@
Figure 3 illustrates the time histories of $x$ and $y$ as well as the planar trajectory of the unicycle. The results clearly demonstrate that the proposed exponentially convergent ESC achieves significantly faster convergence toward the desired equilibrium compared with the traditional ESC method from literature based on first-order Lie bracket design (durr2013lie). Specifically, for the fourth-order objective function considered, the traditional ESC fails to converge within the 100 s simulation window, whereas the proposed design successfully converges within approximately 20 s. Lastly, the proposed exponentially convergent unicycle ESC exhibits a higher radius of oscillation compared to the traditional ESC method. This behavior is due to the control law itself requiring a higher amplitude for the perturbation signal input as observed in grushkovskaya2025extremum; pokhrel2023higher. Techniques to attenuate oscillation such as those in pokhrel2023control; grushkovskaya2024step; yilmaz2025unbiased can possibly resolve this issue and can be applied to the exponentially convergent unicycle ESC. We also note that attenuation of oscillations in the unicycle-based ESC has been successfully tried in robotic experiments as in elgohary2025model.
@@
Figure 3: Simulation results comparing the proposed exponentially convergent unicycle ESC with the traditional ESC based on first-order Lie bracket design from durr2013lie for a fourth-order objective function: (a) $x$ position, (b) $y$ position, and (c) planar trajectory of the proposed unicycle design.
@@ 4.3 Experimental Results
We note here that our design parameters meet the condition of Theorem 1 (see Remark 1). The experimental results are presented in Figures 4 and 5. As shown in Figure 4, the traditional ESC method based on first-order Lie bracket from durr2013lie fails to converge to the true minimum of the objective function even after 1200 s, instead settling near (x,y)=(0.98,1.93)(x,y)=(0.98,-1.93). Due to the fourth-order objective function, the traditional ESC method will not be able to excite high-order Lie brackets and hence it is not possible to guarantee exponential convergence for this approach. Instead, the traditional ESC method will follow a polynomial decay rate, which will be much slower than an exponential decay rate as discussed and clarified in grushkovskaya2025extremum.
In contrast, the exponentially convergent unicycle ESC design converges much faster, first reaching the minimum at approximately 350 s and remaining about the true extremum for the duration of the experiment. The new ESC design clearly outperforms the classic literature ESC design and is validated by its ability to reach the minimum in significantly less time than the traditional approach. It is also important to note that a fourth-order objective function exhibits relatively flat regions near the minimum compared to a second-order objective function. The exponentially convergent unicycle ESC design can successfully steer the robot through these flat regions, as evident in Figure 6. Moreover, the variations in the objective function values become very small when the robot is close to the extremum point, confirming that steady-state convergence has been achieved. It is also important to highlight that the proposed design performed well even with expected resolution issues from the motion capturing system feedback due to the very small values of the fourth-order objective function, especially near the extremum. This is a positive indication about the robustness of the proposed design. The reader is directed to watch the experiment in our YouTube channel (video_known_J).
Figure 4: Experimental results for the traditional ESC design based on first-order Lie bracket from durr2013lie with a fourth-order objective function. (a) $x$ position (b) $y$ position (c) planar trajectory of the robot. Note that the x-position converges to $0.98$ meters (black-dashed), resulting in a 0.02 meter error, and the y-position converges to 1.93-1.93 (black-dashed), which results in a 0.07 meter error. The true minimum is in red.
Figure 5: Experimental results for the exponentially convergent unicycle ESC design with a fourth-order objective function. (a) $x$ position, (b) $y$ position, (c) planar trajectory of robot, (d) objective function, $J$, (e) linear velocity, $\mathrm{v}$. The reader can watch the experiment at our YouTube channel (video_known_J).
@@
The results of this experiment are presented in Figure 7. As shown, the robot successfully reaches and oscillates around the location of maximum light intensity. The light sensor measurements decrease as the robot approaches the source, and the mean value of the signal is minimized toward the end of the experiment, despite small oscillations caused by the robot’s continuous motion. These results further confirm that the proposed unicycle ESC design can autonomously steer the robot through an unknown and spatially varying light field, thereby validating its real-time, model-free source-seeking capability. The reader is directed to watch this experiment video in YouTube (video_light).
@@
Figure 7: Experimental results for light source-seeking with the exponentially convergent ESC model. (a) $x$ position, (b) $y$ position, (c) planar trajectory of robot, (d) objective function, $J$, (e) linear velocity, $\mathrm{v}$. The reader can watch the experiment at our YouTube channel (video_light).
4.4 Response to Noise
In this section, we investigate the robustness of the exponentially convergent unicycle ESC design to noise. To ensure that the numerical simulations capture, as best as possible, the real-world disturbances, we add noise to the measurement of the objective function $J$.
We begin by considering numerical simulations for the exponentially convergent unicycle and include three different levels of noise. Noise is added to the objective function using the Simulink Random Number block with zero mean and varying variance values to achieve different levels of noise. A fixed seed is used to ensure repeatability when using the Random Number block. We conduct numerical simulations using the parameters from Table 1 and provide the results in Figure 8. From these results, the unicycle still converges to the desired position even with noise between $10^{-2}$ and $2\times 10^{-2}$, which is roughly $38\%$ to $77\%$ of the initial value of the objective function. Moreover, we also report robust response to noise; that is, the lesser the intensity of the noise is, the lesser the radius of oscillation in $x$ and $y$.
Figure 8: Simulation results for the exponentially convergent unicycle ESC with three added measurement noise (low, medium and high). (a) $x$ position, (b) $y$ position, (c) objective function, $J$.
For the experiment for the known objective function and the light source-seeking, noise was already present in the system from the MCS and analog light sensor. We provide 100 s of objective function measurements while the robot is stationary at the initial position (i.e., without motion) from the MCS and analog light sensor in Figure 9. Yet, as reported in the previous subsection, all experiments were successful in convergence despite of these inherent noises.
Figure 9: Objective function measurement for the experiment for 100 s with stationary robot at respective initial position in Table 2 and Table 3. (a) Measured objective function from MCS. (b) Measured objective function from the analog light sensor.
4.5 Response to Delay
We investigate the robustness of the exponentially convergent unicycle ESC design to delay. To best relate the numerical simulations to real-world applications, we include a delay in the objective function, $J$.
We consider three different levels of delay for the numerical simulations. The Simulink Transport Delay block, with an initial value equivalent to the initial value of the objective function, is used. We use the parameters from Table 1 and present the results in Figure 10. We observe that even with a delay of up to $0.2$ s, the exponentially convergent unicycle still converges to the desired position quickly.
Figure 10: Simulation results for exponentially convergent unicycle ESC with delay. (a) $x$ position, (b) $y$ position, (c) objective function, $J$.
Furthermore, we note that we did not compensate for delay in the experiment that is inherently present in the system due to the latency between the obtained measurements and the latency in setting the commanded velocity by the robot. As seen in Figure 5 and Figure 7, the robot was able to converge to the desired position regardless of the delays present in the system.
4.6 Response to Different Initial Values in Simulations and Experiment
We investigate the response of the exponentially convergent unicycle ESC with different initial conditions to prove the robustness of the system, and draw some observations about potential gap between simulations and experiments.
We begin by testing three different starting positions in both simulation and experiment. First, we consider the following three sets of initial conditions: (a) x(0),y(0)=0.6,1.251x(0),\ y(0)=0.6,\ -1.251, (b) x(0),y(0)=1.6,2.6x(0),\ y(0)=1.6,\ -2.6, (c) x(0),y(0)=1.7552,1.6126x(0),\ y(0)=1.7552,\ -1.6126. These initial conditions were selected to reflect different positions that are about the same radius from the extremum point to test directional consistency in convergence. The simulations are conducted using the parameters from Table 1, and the experiments are conducted using the parameters from Table 2. The simulation results are provided in Figure 11, and the experimental results are provided in Figure 12. In both simulation and experiments, we show that regardless of the initial position direction, the unicycle or robot can successfully converge to the desired position, demonstrating consistency and robustness of convergence from different directions.
Figure 11: Simulation results for different initial conditions representing the same radius from the extremum to test consistency in convergence of the exponentially convergent unicycle ESC with respect to direction. (a) x(0),y(0)=0.6,1.251x(0),\ y(0)=0.6,\ -1.251, (b) x(0),y(0)=1.6,2.6x(0),\ y(0)=1.6,\ -2.6, (c) x(0),y(0)=1.7552,1.6126x(0),\ y(0)=1.7552,\ -1.6126.
Figure 12: Experimental results for different initial conditions representing the same radius from the extremum to test consistency in convergence of the exponentially convergent unicycle ESC with respect to direction. (a) x(0),y(0)=0.6,1.251x(0),\ y(0)=0.6,\ -1.251, (b) x(0),y(0)=1.6,2.6x(0),\ y(0)=1.6,\ -2.6, (c) x(0),y(0)=1.7552,1.6126x(0),\ y(0)=1.7552,\ -1.6126.
Furthermore, we extend the distance between the initial and desired positions and show, in simulation, the ability of the exponentially convergent unicycle ESC to reach the desired position from a further distance. The initial condition for this simulation is taken as, x(0),y(0)=0.3,0.7x(0),\ y(0)=-0.3,\ -0.7 and the result is shown in Figure 13. The unicycle is still able to converge to the desired position even when starting further away from the desired position.
Figure 13: Simulation result for an initial position that is further away from the desired position with the exponentially convergent unicycle ESC. x(0),y(0)=0.3,0.7x(0),\ y(0)=-0.3,\ -0.7.
In the experiment, due to the hardware limitations of the TurtleBot3 robot, we were not able to test initial conditions that were further away from the desired position.
Overall, the exponentially convergent unicycle ESC does not show dependence on the initial position directions with respect to the extremum in both simulations and experiments. However, a gap between theory/simulations on the one hand and experiments on the other, was observed as the exponentially convergent unicycle has hardware limitations preventing it from converging if the initial condition is set very far from the extremum; a solution for that could be a future work that develops a bounded update rate design upgrade for the exponentially convergent ESC, analogous to the bounded update design upgrade in scheinker2017bounded with respect to the traditional ESC design in durr2013lie.
4.7 Comparison between Simulation and Experiment Results
We further investigate a direct comparison between the simulation results and the experimental results to evaluate if the performance of the exponentially convergent unicycle ESC matches the theoretically predicted exponential convergence rate provided in Section 3 that is illustrated by simulations.
To begin, we note that the unicycle dynamics model does not perfectly capture the true dynamics of the TurtleBot3 robot. Inertial terms are not captured in the unicycle dynamics as the unicycle is treated as a point mass. Also, dissipation terms, such as friction, are not captured within the unicycle dynamics. For these reasons, it is not possible to directly equate the simulation and experiment parameters. That is, Table 1 and Table 2 are not aligned in terms of the parameters, partly due to the differences in the unicycle model and the real-world system of TurtleBot3. To compare the simulation results with its theoretically predicted behavior vs. the experimental result, we aimed to match the initial convergence rate for the states between the two systems (simulation-based and experiment-based) in a similar fashion to what could be found in ESC literature when one compares the convergence rate of two different ESC systems, see for example the comparisons in elgohary2025extremum. Analyzing Table 1 and Table 2, we see that at $t=0$ s, $\dot{y}(0)$ is the same for both systems and the only difference between the two systems is in $\dot{x}(0)$ with the simulation-based system possessing a much higher initial convergence rate in $\dot{x}$ compared to the experiment-based system. That is, for the experiment-based system, $\dot{x}=0.2199$, and for the simulation-based system, $\dot{x}(0)=705.724$. We now introduce a new set of parameters, in Table 4, to reduce the initial convergence rate of $\dot{x}$ in the simulation-based system as much as possible to allow for a realtively fairer comparison with the experiment-based system. The experimental parameters are the same as in Table 2. With the new parameters in Table 4 for the simulation-based system, the initial convergence rate for $\dot{x}$ becomes $\dot{x}(0)=5.5382$. Now, the results that compare the experiment-based system and the simulation-based system that theoretically is guaranteed to have exponential convergence rate are provided in Figure 14. We observe that the simulation-based system, even with relatively higher initial convergence rate in $\dot{x}$ (i.e., more advantageous initial speed), is slower than the experimental-based system. This shows that the proposed unicycle ESC follows the theoretical predictions in Section 3 and their simulations; however, one need to keep in mind the reported limitation in Subsection 4.6 where the real-world system may not converge for initial conditions that are very far from the extremum.
Table 4: Simulation Parameters for Lower $\dot{x}$ Convergence Rate
Parameter Value
$C_{1},\ C_{2}$ 1, 11,\ 1
$a$ $0.1$
$c$ $0.5$
$\varepsilon$ $0.075$
$\Omega$ 1.4rad/s1.4\ $\text${rad/s}
x(0),y(0)x(0),\ y(0) 1.6,1.4m1.6,\ -1.4\ $\text${m}
$x_{d},\ y_{d}$ 1,2m1,\ -2\ $\text${m}
Figure 14: Comparison between experiment-based system (blue) and simulation-based system (red) with close, but still higher initial convergence rate demonstrates that the proposed exponentially convergent unicycle design follows the predicted theoretical behavior.
We also note that for the model-free light source-seeking experiment, we use different parameters compared to the known objective function experiment and simulations. This is in part because the light distribution is not known, and the scale of the measurements received from the light sensor differs from the cases when the objective function is known as fourth-order polynomial, hence differences in tuning are expected.
@@ 5 Conclusion and future work
This paper provided a novel, first-of-its-kind exponentially convergent unicycle design inspired by significant recent results in higher-order Lie bracket approximations (grushkovskaya2025extremum; pokhrel2023higher). We proved exponential stability of the proposed design for objective functions that behave locally as a fourth-degree polynomial. Additionally, we validated the results by simulations and, more importantly, via experiments. In simulations, we showed the effective response of the proposed design under different initial conditions and under added levels of measurement delays/noise. In our experiments, we did not only validate the ability of the proposed design to operate with known objective functions and with different initial conditions, but we also validated its ability to operate in a complete model-free condition in a light source-seeking experiment. Furthermore, we compared for a case study the experimental results behavior vs. a simulation result behavior to confirm that the experimental result follow the predicted theoretical behavior.